Grade 11Algebra

how to solve the question The GM of the roots of the equation x^2-18x+9=0

Sarthak Brahma
10 Years agoGrade 11

1 Answer

Riddhish Bhalodia
ApprovedApproved Tutor Answer10 Years ago
Let the roots be A and B, then GM of A and B is

GM = \sqrt{AB}
Now, so what we just need is the product of roots AB
which by quadratic equations is equal to
AB = 9/1 = 9
hence GM = 3